OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (0, 0) and (3, 30).

OpenStudy (helder_edwin):

if u have the points \((a,b)\) and \((c,d)\) then the slope is \[ \large m=\frac{d-b}{c-a}=\frac{b-d}{a-c} \]